# Post-detection inference for sequential changepoint localization

### v2

@@ 1 Introduction
We consider the following problem of sequential change analysis in a general space $\mathcal{X}$: a sequence of $\mathcal{X}$-valued observations $X_{1},X_{2},\cdots$ arrives sequentially, such that for some unknown time $T\in\mathbb{N}\cup\{\infty\}$ (the changepoint),
$X_{1},X_{2},\cdots,X_{T-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1},$ (1.1)
for some pre-change distribution $F_{0}\in\mathcal{P}$ and some post-change distribution $F_{1}\in\mathcal{P}$, where $\mathcal{P}$ denotes a class of probability distributions on $\mathcal{X}.$ We start by assuming $F_{0}$ and $F_{1}$ are known, and generalize our methods to encompass settings where $F_{0}$ and $F_{1}$ are only known to lie in some classes, but are not known exactly. A sequential changepoint detection algorithm $\mathcal{A}$ is simply a data-dependent stopping rule $\tau$ to raise an alarm that a change has previously occurred. Our aim in this paper is to construct confidence sets for the unknown changepoint $T$ after stopping at $\tau$, assuming only black-box access to $\mathcal{A}$ (and ideally no further assumptions or restrictions). In other words, we hope to obtain a set $\mathcal{C}\subseteq\{1,\dots,\tau\}$ such that
$\mathbb{P}_{T}(T\in\mathcal{C})\geq 1-\alpha,$
where $\mathbb{P}_{T}$ is the data distribution in (1.1), i.e. having the true changepoint at $T$. We will show for algorithms $\mathcal{A}$ with a bounded average run length (ARL), no set $\mathcal{C}$ can satisfy such a guarantee, but if $\mathcal{A}$ has probability of false alarm (PFA) bounded above by $\alpha$, such a guarantee is achievable. In either case, another natural criterion for $\mathcal{C}$ is for it to satisfy
$\mathbb{P}_{T}(T\in\mathcal{C}|\tau\geq T)\geq 1-\alpha,$
meaning that $\mathcal{C}$ only needs to cover if algorithm $\mathcal{A}$ only stops after the true changepoint $T$. We show how to construct such sets $\mathcal{C}$ for both types of algorithms (bounded ARL or PFA).
Sequential (also known as the online or quickest) changepoint detection is a well-studied field, where the goal is to detect that a change has occurred as soon as possible, is crucial in applications requiring timely responses to distributional shifts, such as quality control, cybersecurity, medical diagnostics, and so on. The field traces back to early work on statistical process control. Shewhart [19] introduced the concept of a control chart, which computes a test statistic for each sample, with values within a threshold indicating the process remains in control. Page [16] proposed the CUSUM (cumulative sum) procedure, which computes the cumulative log-likelihood ratio under the assumption of two exactly known distributions. Another influential approach is the Shiryaev-Roberts (SR) procedure [21, 18]. These early methods, however, assume the pre- and post-change distributions to be fully specified. To relax the assumption of known post-change distributions, the Generalized Likelihood Ratio (GLR) procedure was introduced [15, 22]. Tartakovsky et al. [23] provide comprehensive overviews of these classical methods, including CUSUM, SR, and GLR, and their numerous variants. Several nonparametric change-point detection algorithms have been developed, such as kernel-based methods [7, 14] and e-detectors [20].
Localizing the changepoint is another fundamental and crucial problem. As far as we are aware, the existing approaches to changepoint localization focus on offline or batch settings, where a fixed amount of data is available at once, and they do not generalize to analysis at data-dependent stopping times. Hinkley [8] shows some asymptotic results about the maximum likelihood estimator (MLE) of the changepoint and also provides a confidence set based on likelihood ratio statistics. Yao [26] provides an approximation of the distribution of the MLE of the changepoint. Darkhovskh [3] presents a nonparametric changepoint estimator based on Mann-Whitney statistics. Carlstein [1] proposes a class of strongly consistent estimators for the changepoint in the nonparametric setting. Worsley [25] develops likelihood ratio-based confidence regions and tests for a changepoint for exponential family random variables. Dümbgen [5] constructs nonparametric confidence regions for changepoint by testing for each candidate time-point $t$ at level $\alpha$ using bootstrapping, achieving asymptotic coverage of $1-\alpha$, as the sample size $n\to\infty$. A more recent method [10] constructs confidence sets by evaluating a local test statistic on a triplet of time points and combining the results of the local tests with a weighted Bonferroni correction. To re-emphasize, all these methods are for the offline setting, while our main focus is on the sequential problem of constructing confidence sets after stopping, which has been largely overlooked in the literature. This gap is significant because in most real-world applications, data arrives sequentially, and making timely inferences about the changepoint immediately upon detecting a change is essential. Thus, developing methodologies for post-detection inference in a sequential framework is a significant step forward in changepoint analysis.
The key idea behind constructing the confidence interval is simple: for each candidate $t\in\{1,\cdots,\tau\}$, we will test whether $t$ is the true changepoint and form the confidence set by inverting these tests. Our approach is conceptually similar to [5], it adapts those ideas to the sequential setting in a nontrivial manner. For example, we show that for change detection algorithms that have finite average run length (ARL), it is impossible to obtain finite length confidence intervals with marginal coverage guarantees. Further, for $1-\alpha$ conditional coverage, we show that we need to test the $t$-th hypothesis at level $\alpha\times\mathbb{P}_{\infty}(\tau\geq t)$ (a quantity we do not know apriori), whereas testing for each $t$ at level $\alpha$ works in the offline setting. We develop simulation based approaches to get around these hurdles.
There are two settings for the changepoint problem: non-partitioned and partitioned. The non-partitioned changepoint problem assumes that the data-generating distribution changes from one unknown distribution $F_{0}\in\mathcal{P}$ to another unknown distribution $F_{1}\in\mathcal{P}$, without requiring any pre-specified partitioning of $\mathcal{P}$ into pre- and post-change distribution classes (eg: detecting a change in a Gaussian mean from any value to any other value). In contrast, the partitioned setting assumes that we know some partition of $\mathcal{P}$ into $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ such that $F_{0}$ and $F_{1}$ are known to lie in $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ respectively. This formulation is useful when changes within $\mathcal{P}_{0}$ are not of interest, and we only seek to detect shifts from $\mathcal{P}_{0}$ to $\mathcal{P}_{1}$ (eg: detecting a change in Gaussian mean from negative to positive). Our method is designed for the partitioned setting. Extensions to the non-partitioned case are left for future work.
1.1 Our contributions
Here we summarize the key contributions of the paper.
  • For detection algorithms that control PFA, we propose Algorithm 1 for constructing a confidence set for the changepoint and establish its unconditional and conditional coverage (Theorem 2.2, Corollary 2.6), assuming known pre- and post-change distributions.
  • For detection algorithms with bounded ARL, we show in Proposition 2.4 that we cannot have an unconditional coverage guarantee and propose Algorithm 2, which is a modification of Algorithm 1. We provide its conditional coverage guarantee in Theorem 2.7.
  • We extend our framework to composite pre- and post-change classes in Algorithm 3 and Algorithm 4 and provide the coverage guarantees in Theorem 3.1 and Theorem 4.2.
  • We demonstrate that the approach is not overly conservative (and thus practical) by examining the achieved coverage and size of constructed confidence sets in extensive simulations.
We point out again that for obtaining the aforementioned guarantees, we only need to have black-box access to the detection algorithm $\mathcal{A}$. It need not have formal guarantees, we do not need to know its inner workings, and just need to be able to run $\mathcal{A}$ on any stream of data.
1.2 Paper outline
In Section 2, we construct our confidence sets for the changepoint for known pre- and post-change and study its unconditional and conditional coverage probabilities. Section 3 extend this approach to handle composite post-change, assuming the pre-change distribution to be known. In Section 4, we finally extend this approach to handle composite pre- and post-change setting. Section 5.1 and Section 5.2 present a comprehensive set of simulation studies that validate our theoretical findings and demonstrate the practical efficacy of our approach. This article is concluded in Section 6. Omitted proofs can be found in Appendix A. An additional illustrative example using the Poisson distribution is given in Appendix B. Appendix B. Some additional experimental results are provided in Appendix C and some heuristic algorithms are presented in Appendices D and E, which are computationally more efficient. A summary of comparisons of the heuristic and non-heuristic methods is provided in Appendix E, pointing at open avenues for further theoretical work.
2 Known pre- and post-change distributions
We now address the question of constructing confidence sets for the changepoint, assuming that our change detection procedure stopped after a changepoint. We start from the simplest case when both pre- and post-change distributions are known, addressing more general cases in later sections.
[…]

### v3

@@ 1 Introduction
Consider the following general problem of sequential change analysis in an arbitrary space $\mathcal{X}$ (which could be the Euclidean space $\mathbb{R}^{d}$, or images, or anything else). A sequence of $\mathcal{X}$-valued observations ${\bf X}=X_{1},X_{2},\cdots$ arrives sequentially, such that for some unknown time ${}T\in\mathbb{N}\cup\{\infty\}$ (the changepoint),
$X_{1},X_{2},\cdots,X_{T-1}\stackrel{{\scriptstyle}}{{\sim}}F_{0}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle}}{{\sim}}F_{1},$ (1.1)
for some $F_{0}$ in a pre-change class $\mathcal{P}_{0}$ and $F_{1}$ in a post-change class $\mathcal{P}_{1}$, where $\mathcal{P}_{0}$ and $\mathcal{P}_{0}$ denote some non-intersecting classes of probability distributions. Initially, we shall assume all observations to be independent, although our methodology later extends to handling dependent data as well.
Suppose the user has already decided to use a sequential changepoint detection algorithm $\mathcal{A}$ for this task, which is a mapping from the data sequence $\bf X$ to a stopping time [^1] While sequential change detection methods are commonly described solely by stopping times, we describe them by algorithms $\mathcal{A}$ because our method will feed in simulated (synthetically generated) data sequences $\bf X^{\prime}$ into $\mathcal{A}$; we find it easier to describe our methodology when making $\mathcal{A}$ explicit. $\tau=\mathcal{A}({\bf X})$ at which it raises an alarm that a change may have previously occurred. When $\mathcal{A}$ raises an alarm at time $\tau$, we do not know if it is a false alarm, and if it is not, then we only know that there has been some possible change before $\tau$, but not when that change occurred.
Our aim here is to construct confidence sets for the unknown changepoint $T$ after stopping at $\tau$, assuming only black-box access to $\mathcal{A}$ (and ideally no further assumptions or restrictions). By “black-box”, we mean that we can run $\mathcal{A}$ on any input data sequences we like, but we do not require any further insight into its inner workings. The fact that we do not require knowledge of internal details of $\mathcal{A}$ is a particularly important feature; it implies that our changepoint localization method can wrap around any existing workflow for changepoint detection — it could be employed both with simple classical statistical algorithms (like CUSUM), and also with complex modern pipelines in the information technology industry that may involve heuristic machine learning algorithms (to deal with complex data types).
As an initial goal, one may hope to obtain a set $\mathcal{C}\subseteq\{1,\dots,\tau\}$ such that
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C})\geq 1-\alpha,$ (1.2)
where $\mathbb{P}_{F_{0},T,F_{1}}$ is the data distribution in (1.1), i.e. having the true changepoint at $T$. But this goal is harder to achieve than may initally appear; we elaborate below.
There are two classes of algorithms $\mathcal{A}$ in the literature: (i) algorithms with probability of false alarm (PFA) bounded by $\delta<1$; we call these “bounded PFA” for short, and (ii) algorithms with PFA equal to one, but have a bounded average run length to false alarm (ARL2FA, or just ARL for short); we call these “bounded ARL”. Since algorithms with bounded ARL have PFA equal to one, false detections occur with a nontrivial frequency. But the flipside is that bounded ARL algorithms typically have smaller detection delay than bounded PFA algorithms. Typically, detection delays of bounded ARL algorithms can be bounded by quantities independent of the changepoint $T$, but this guarantee is impossible to achieve with bounded PFA algorithms. In this paper, we remain agnostic to the choice of $\mathcal{A}$.
For algorithms with a bounded ARL, we prove that no set $\mathcal{C}$ can satisfy guarantee (1.2) (for algorithms with bounded PFA, we show how to achieve (1.2) in the Supplementary). This impossibility result (which we prove formally later) intuitively holds because it is meaningless to ask for coverage if the algorithm stopped before the changepoint; indeed, on the event $\{\tau<T\}$, we have no information to rule out the possibility that $T=\infty$, and thus the only way to cover $T$ is to output the set of all natural numbers $\mathbb{N}$. Thus, an arguably more sensible goal is to have
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,$ (1.3)
meaning that $\mathcal{C}$ only needs to cover if algorithm $\mathcal{A}$ only stops after the true changepoint $T$. This definition has been considered by, for example, [5] and [39, 36], who achieved it (in a certain asymptotic sense) in much more restricted settings than the current investigation.
The aforesaid impossibility result no longer applies to the new goal (1.3). It is important to note that we do not prescribe which $\mathcal{A}$ to use in which situation, but simply enable post-hoc inference on $T$ for any user-chosen $\mathcal{A}$. One can view our techniques as providing a wrapper to convert $\mathcal{A}$ from producing only a stopping time $\tau$ to also producing a confidence set for $T$  [^2] In some sense, our goal can be viewed in an analogous way to conformal inference, which wraps around an arbitrary prediction algorithm to convert a “point prediction” into a prediction set, by running the same algorithm on many hypothetical datasets (for full conformal prediction). However, our methods are not distribution-free, and rely on smartly using simulations to accomplish the goal.. As we survey below, there is a rich literature on designing change detection algorithms $\mathcal{A}$, but a substantial paucity of inference methodology for localizing the change after detection; our focus is on the latter goal, assuming the user has chosen a suitable algorithm $\mathcal{A}$ for detection.
While inference for $T$ is intrinsically interesting and important, a side benefit of confidence sets for $T$ is that one can use these to derive post-hoc confidence sets for unknown parameters of $F_{0}$ and $F_{1}$. To clarify, suppose $\mathcal{P}_{i}=\{F_{\theta}\}_{\theta\in\Theta_{i}};i=0,1$. Our method that delivers (1.3) also delivers confidence sets $\mathcal{C}^{\prime}\subseteq\Theta_{0}\times\Theta_{1}$ such that $\mathbb{P}_{\theta_{0},T,\theta_{1}}((\theta_{0},\theta_{1})\in\mathcal{C}^{\prime}\mid\tau\geq T)\geq 1-\alpha,\forall\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1},$ where $\mathbb{P}_{\theta_{0},T,\theta_{1}}$ is the distribution of the sequence $X_{1},\dots,X_{T-1}\stackrel{{\scriptstyle}}{{\sim}}F_{\theta_{0}}$, $X_{T},X_{T+1},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{\theta_{1}}$. We stress that our confidence sets do not rely on asymptotics and are valid in finite samples.
1.1 Related work
Prior work on post-detection confidence sets for $T$. The problem of constructing confidence sets for the changepoint after stopping has been largely overlooked in the literature. This gap is significant because in most real-world applications, data arrives sequentially, and it can be essential to make timely inferences about the changepoint immediately upon detecting a change.
[…]
In sharp contrast, our proposed confidence sets for $T$ are very general: We make no assumptions on the detection procedure $\mathcal{A}$, and our framework acts as a versatile wrapper that can be layered on top of any sequential change detection algorithm. It is not restricted to exponential families — it works for any known $F_{0}$ and any nonparametric post-change class $\mathcal{P}_{1}$ on any general space (e.g., $\mathbb{R}^{d}$) and offers non-asymptotic coverage guarantees. It also extends to composite pre-change classes under some assumptions. Thus, our framework is the first nonparametric, theoretically sound, and widely applicable practical tool for sequential change localization tasks. We summarize the comparison in Table 1.1.
Table 1.1: Comparison of existing methods and our proposed approach
Pre-, Post- Existing Methods [5, 36] Our Proposed Method
change distributions Assumptions Guarantee Assumptions Guarantee
Known, Known $\mathcal{A}=$ CUSUM. $F_{0},F_{1}$ Asymptotic No assumption Nonasymptotic
[…]
Known, Unknown No existing work No assumption Nonasymptotic
($F_{0},\mathcal{P}_{1}$) (any $\mathcal{A},F_{0},\mathcal{P}_{1}$) (Theorem 4.4)
Unknown, Unknown No existing work Assumption 3.3 Nonasymptotic
($\mathcal{P}_{0},\mathcal{P}_{1}$) on $\mathcal{A},\mathcal{P}_{0}$ (any $\mathcal{P}_{1}$) (Theorem 4.5)
Literature on sequential change detection. In sequential change detection, the goal is to detect that a change has occurred as soon as possible, but not to provide inferential tools for when the change occurred. This methodology has proved crucial in many applications, such as quality control, cybersecurity, medical diagnostics, and so on. Classical methods such as Shewhart control charts [25], the CUSUM procedure [18], and the Shiryaev-Roberts approach [27, 22] assume known pre- and post-change distributions. To address scenarios with unknown post-change distributions, generalized likelihood ratio (GLR) procedures [17, 28] were developed. Comprehensive reviews of these methods are provided by [30]. More recently, kernel-based techniques [9], e-detectors [26], etc., have emerged to handle nonparametric settings.
Offline vs. online change localization. As far as we are aware, most existing approaches to changepoint localization focus on offline settings, where a fixed amount of data is available at once, and they do not generalize to analysis at data-dependent stopping times. Classical works, e.g., [10] study the asymptotic behavior of the MLE for the changepoint and construct likelihood-ratio-based confidence sets. [3] proposed some nonparametric estimators. [35] developed likelihood ratio-based confidence sets for exponential family distributions, while [6] used bootstrap-based tests for each candidate time to construct nonparametric confidence sets with asymptotic coverage. Recently, [31] studies optimal rates for both changepoint detection and localization. [13] constructs confidence sets by evaluating a local test statistic. To re-emphasize, all these methods are for the offline setting. In the sequential setting, [29, 7, 37, 1] develop point estimators for $T$; but very few works construct confidence sets [5, 36], whose drawbacks were discussed earlier.
Inference on $\theta_{0},\theta_{1}$ after stopping. As with the problem of post-detection inference on the changepoint $T$, there is limited literature on valid estimation or inference for the pre- and post-change parameters $\theta_{0}$ and $\theta_{1}$ post-stopping. [39] constructs asymptotic bias-corrected point estimates and confidence intervals for the post-change mean after a mean change is detected by a CUSUM procedure. [38] generalizes these results for the case when there is a possible change in the variance as well. These methods are also discussed in [36]. However, these approaches share several limitations with the existing post-detection changepoint inference methods, as mentioned earlier: they rely on asymptotics, assume specific parametric forms, and apply only to the CUSUM procedure.
1.2 Contributions and paper outline
  • We propose a general framework for constructing a confidence set for the changepoint based on inverting tests obtained using a universal threshold, assuming all observations to be independent. We show that it provides $1-\alpha$ conditional coverage for known pre-change and arbitrary (nonparametric) post-change classes under no additional assumption in Theorem 3.2. We also extend it for composite pre-change under a suitable assumption in Theorem 3.5.
  • We propose an alternative method of constructing the tests using adaptive thresholds (which is often tighter than the universal threshold), assuming pre- and post-change distributions to be known, in Algorithm 1. We extend this framework to handle parametric pre- and/or post-change classes in Algorithms 2 and 3. These adaptive methods allow the observations to be dependent. We establish corresponding coverage guarantees in Theorems 4.1, 4.4 and 4.5.
  • As a byproduct of our methods, we derive confidence sets for the unknown pre- and post-change parameters and show their coverage guarantees in Theorem 5.1.
In short, we present the first nonparametric, theoretically grounded, and widely applicable post-detection confidence sets that are agnostic to the detection method.
The rest of the paper is organized as follows. In Section 2, we establish some fundamental results showing the impossibility of unconditional coverage and linking conditional confidence sets to hypothesis tests. Section 3 develops a general framework for constructing a confidence set for $T$ based on inverting tests obtained using a universal threshold. In Section 4, we introduce a simulation-based approach for constructing confidence sets using an adaptive threshold in the parametric setting and study their conditional coverage probabilities. In Section 5, we derive confidence sets for the changepoint, as well as the pre- and post-change parameters. Section 6 presents an extensive simulation studies that validate our theoretical findings and demonstrate the practical efficacy of our approach. This article is concluded in Section 7. All the mathematical proofs, implementation details of some algorithms in the paper, and additional experiments are presented in the Supplementary Material.
2 Some Fundamental Results
In this section, we highlight some key observations that shape our methodology of post-detection inference for the change-point $T$. We begin by clarifying some notational conventions that will be used throughout the paper.
Notation. For $t\in\mathbb{N}$, let $\mathbb{P}_{F_{0},t,F_{1}}$ denote joint distribution of the sequence $\{X_{n}\}_{n}$ when $X_{1},\dots,X_{t-1}\stackrel{{\scriptstyle}}{{\sim}}F_{0}$ and $X_{t},X_{t+1},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{1}$. 333 To understand the key ideas in the paper, it will suffice for the reader to focus on the case when all $X_{i}$’s are all independent, and $F_{0}$ and $F_{1}$ denote the pre- and post-change marginal distributions. However, our framework in Section 4 can actually handle dependent data. Under dependence, $F_{0}$ and $F_{1}$ denotes the joint distributions of $X_{1},\dots,X_{t-1}$ and $X_{t},X_{t+1},\dots$ respectively, and is in fact ambivalent to whether $X_{t}$ is drawn independently of $X_{1},\dots,X_{t-1}$ or conditional on it, so we do not emphasize the distinction in our notation. As a corner cases, let $\mathbb{P}_{F_{0},\infty}$ be the joint distribution when $X_{1},X_{2},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{0}$. In parametric settings, we write $\mathbb{P}_{\theta_{0},t,\theta_{1}}$ and $\mathbb{P}_{\theta_{0},\infty}$ as shorthand for $\mathbb{P}_{F_{\theta_{0}},t,F_{\theta_{1}}}$ and $\mathbb{P}_{F_{\theta_{0}},\infty}$ respectively.
It is impossible for any set $\mathcal{C}^{*}\subseteq\{1,\cdots,\tau\}$, to produce an unconditional coverage guarantee at a predetermined level $1-\alpha$, if the detection algorithm has finite ARL (i.e., $\tau$ is finite almost surely even if there is no change). The intuitive reason is that if $\tau$ is finite almost surely, it is possibly a false detection, i.e., $\tau<T$, in which case the set $\{1,\cdots,\tau\}$ itself does not cover $T$. The next proposition formalizes this fact.
Proposition 2.1.
If, for some $\alpha\in(0,1)$, a confidence set $\mathcal{C}^{*}$ that is a subset of $\{1,\cdots,\tau\}$ satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}^{*})\geq 1-\alpha$, $\forall T\in\mathbb{N}$, then $\mathbb{P}_{F_{0},\infty}(\tau=\infty)\geq 1-\alpha$ (meaning that the PFA of the algorithm is at most $\alpha$ and thus the ARL is infinite). Said differently, for any algorithm with finite ARL, it is impossible to construct a $1-\alpha$ confidence set that is only a subset of $\{1,\dots,\tau\}$.
Thus, for post-detection inference, the usual notion of coverage is not an appropriate one. Instead, we focus on conditional coverage (1.3), a more reasonable criterion, which was also considered in prior work [5, 36].       The key idea behind constructing our confidence intervals is simple, but several details are subtle. For each candidate $t\in\{1,\cdots,\tau\}$, we test whether $t$ is the true changepoint from some pre-change class $\mathcal{P}_{0}$ to some post-change class $\mathcal{P}_{1}$:
$H_{0,t}:\;X_{1},\dots,X_{t-1}\sim F_{0}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ X_{t},X_{t+1},\dots\sim F_{1},\text{ for some }F_{i}\in\mathcal{P}_{i};i=0,1.$ (2.1)
Let $\phi_{\alpha^{*}}^{(t)}=\mathds{1}(H_{0,t}\text{ is rejected})$ denote some test for the null $H_{0,t}$ such that
$\mathbb{P}_{F_{0},t,F_{1}}(\phi_{\alpha^{*}}^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha^{*},\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}$ (2.2)
and form the confidence set by inverting these tests:
$\mathcal{C}:=\{t\in\mathbb{N}:t\leq\tau,H_{0,t}\text{ is not rejected}\}$ (2.3)
[…]
Proposition 2.2.
Let $\tau$ be a stopping time and $\mathcal{C}\subseteq\{1,\dots,\tau\}$ be a confidence set satisfying (2.5). Then, the test defined by $\phi^{(t)}(X_{1},\cdots,X_{\tau})=\mathds{1}(t\notin\mathcal{C})$ satisfies
$\mathbb{P}_{F_{0},t,F_{1}}(\phi^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq t),\forall t\in\mathbb{N},F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$ (2.4)
Conversely, given any test $\phi^{(t)}$ for $H_{0,t}$ using data up to $\tau$ satisfying (2.4), the inverted confidence set $\mathcal{C}=\{t\in\mathbb{N},t\leq\tau:\phi^{(t)}(X_{1},\cdots,X_{\tau})=0\}$ satisfies
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\leavevmode\nobreak\ \forall T\in\mathbb{N},F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$ (2.5)
The above result establishes the duality between hypothesis tests and a post-detection confidence set. In the next section, we propose a general framework for constructing confidence sets for $T$ by inverting tests based on the above.
3 A general framework for nonparametric post-change
Let $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ denote the pre- and post-change classes respectively, $\mathcal{P}_{0}\cap\mathcal{P}_{1}=\emptyset$. Our goal is to construct post-detection confidence sets for the changepoint $T$ by inverting tests for $H_{0,t}$, as formalized in Proposition 2.2.
To motivate the general construction, we begin with the simplest case when both pre- and post-change distributions $F_{0}$ and $F_{1}$ (with densities $f_{0}$ and $f_{1}$) are known. Then, the likelihood of the data $X_{1},\cdots,X_{T-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0}$, $X_{T},\cdots,X_{n}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1}$ is given by $L_{T}^{n}:=\prod_{i=1}^{T-1}f_{0}(X_{i})\prod_{i=T}^{n}f_{1}(X_{i})$, and the maximum likelihood estimator of $T$ [10] can be written as $\hat{T}_{n}=\operatorname*{arg\,max}_{1\leq j\leq n}\prod_{i=j}^{n}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}.$ In our setting, having data is observed up to a stopping time $\tau$, the analogous estimator is obtained by replacing $n$ with $\tau$ [37, 36]:
Now, a prototypical likelihood-ratio-based test statistic is the following: $M_{t}=\displaystyle\max_{1\leq i\leq\tau}{L_{i}^{\tau}}/{L_{t}^{\tau}},\text{ if }t\leq\tau<\infty$ and $M_{t}=-\infty\text{ if }t>\tau\text{ or }\tau=\infty$. It simplifies to
$M_{t}=\begin{cases}\prod_{i=t}^{\hat{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})},&\text{ if }t<\hat{T}\leq\tau<\infty\\ 1,&\text{ if }t=\hat{T}\leq\tau<\infty\\ \prod_{i=\hat{T}}^{t-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})},&\text{ if }\hat{T}<t\leq\tau<\infty\\ -\infty,&\text{ if }t>\tau\text{ or }\tau=\infty.\end{cases}$ (3.2)
One natural way to extend this test statistic for the general nonparametric setting is to replace the likelihood ratio processes by e-processes [19, 8], which have emerged as a central tool in anytime-valid inference. While e-processes have been used for sequential testing [33, 20], estimation [34], and change detection [26], we employ them for the first time for changepoint localization. For $t\in\mathbb{N}$, we define two different types of e-processes:
  • Forward $t$-delay e-process under $\mathcal{P}_{1}$ (against $\mathcal{P}_{0}$). This is a stochastic process $\{R^{(t)}_{n}\}_{n\geq t}$ adapted to the forward filtration $\{\mathcal{F}^{(t)}_{n}\}_{n=t,t+1,\cdots}$, where $\mathcal{F}^{(t)}_{n}:=\sigma(X_{t},X_{t+1},\dots,X_{n}),$ such that for all stopping times $\kappa$ with respect to $\{\mathcal{F}^{(t)}_{n}\}_{n\geq t}$, $\sup_{F_{1}\in\mathcal{P}_{1}}\mathbb{E}_{F_{1}}[R^{(t)}_{\kappa}]\leq 1.$ Ideally, this process should have power against $\mathcal{P}_{0}$, i.e., $R^{(t)}_{n}$ should grow to infinity (as $n\to\infty$) under $\mathcal{P}_{1}$.
  • Backward $t$-delay e-process under $\mathcal{P}_{0}$ (against $\mathcal{P}_{1}$). This is a stochastic process $\{S^{(t)}_{n}\}_{n<t}$ adapted to the backward filtration $\{\mathcal{G}^{(t)}_{n}\}_{n=t-1,t-2,\cdots}$, where $\mathcal{G}^{(t)}_{n}:=\sigma(X_{n},X_{n+1},\cdots,X_{t-1}),$ such that for all stopping times $\kappa$ with respect to $\{\mathcal{G}^{(t)}_{n}\}_{n<t}$, $\sup_{F_{0}\in\mathcal{P}_{0}}\mathbb{E}_{F_{0}}[S^{(t)}_{\kappa}]\leq 1.$ Ideally, this process should have power against $\mathcal{P}_{1}$, i.e., $S^{(t)}_{n}$ should grow to infinity (as $n\to\infty$) under $\mathcal{P}_{0}$.
Such powered e-processes for $\mathcal{P}_{1}$ against $\mathcal{P}_{0}$ (and vice versa) exist if the total variation distance between the two convex hulls of $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ is positive (see e.g., Section 3.4 of [21]). We require all $X_{i}$’s to be independent for both powered (nontrivial) e-processes to be adapted to their respective filtrations.
The final ingredient we need is $\hat{T}$, which is a preliminary point estimate of $T$.  [^4] Here, we again note that our confidence set based on a point estimate $\hat{T}$ can be viewed analogously to the conformal prediction set, which is constructed around a preliminary “point prediction”. For example, one can construct such an estimator based on e-processes under $\mathcal{P}_{0}$. To elaborate, let $\{E^{(t)}_{n}\}_{n\geq t}$ be some forward $t$-delay e-process under $\mathcal{P}_{0}$ (against $\mathcal{P}_{1}$). Then, one reasonable estimate is
$\hat{T}=\operatorname*{arg\,max}_{1\leq t\leq\tau}E^{(t)}_{\tau}.$ (3.3)
This can be viewed as a nonparametric generalization of the maximum likelihood estimator for known pre- and post-change defined in (3.1).
With the above ingredients in place, we construct a test statistic for $H_{0,t}$ by comparing it against a competing hypothesis $H_{0,T}$. Suppose $t<T$. Then under $H_{0,t}$ (which is false), the segment $X_{t},\dots,X_{T-1}$ follows a post-change distribution from $\mathcal{P}_{1}$ (because under $H_{0,t}$, the change occurred at $t$), whereas under $H_{0,T}$ (which is true), it follows a pre-change distribution from $\mathcal{P}_{0}$. Outside this segment $X_{t},\dots,X_{T-1}$, both hypotheses agree on the distribution of the observations, and hence those are non-informative for testing between $H_{0,t}$ and $H_{0,T}$. Therefore, we use a forward e-process $R^{(t)}_{T-1}$ that accumulates evidence in favor of $\mathcal{P}_{1}$ against $\mathcal{P}_{0}$ over this segment. Similarly, if $t>T$, we employ a backward e-process $S^{(t)}_{T}$ to accumulate evidence for $\mathcal{P}_{0}$ against $\mathcal{P}_{1}$. Since the true change-point $T$ is unknown, we replace it with an estimated change-point $\hat{T}$. This naturally leads us to define the following:
$M_{t}=\begin{cases}R^{(t)}_{\hat{T}-1},&\text{ if }t<\hat{T}\leq\tau<\infty\\ 1,&\text{ if }t=\hat{T}\leq\tau<\infty\\ S^{(t)}_{\hat{T}},&\text{ if }\hat{T}<t\leq\tau<\infty\\ -\infty,&\text{ if }t>\tau\text{ or }\tau=\infty.\end{cases}$ (3.4)
[…]
Theorem 3.1.
$\mathbb{P}_{F_{0},t,F_{1}}[M_{t}\geq 2/\alpha]\leq\alpha,$ for any $t\in\mathbb{N},F_{0}\in\mathcal{P}_{0},$ and $F_{1}\in\mathcal{P}_{1}.$
While this bound may be conservative, it is remarkably universal: it holds without any assumptions on the distribution classes $\mathcal{P}_{0}$, $\mathcal{P}_{1}$, or on the detection method $\mathcal{A}$. For constructing a valid confidence set, Proposition 2.2 implies that we require tests of level $\alpha\cdot\mathbb{P}_{F_{0},\infty}(\tau\geq t)$, whose value is of course not known to us. To circumvent this issue, we first assume that the pre-change distribution $F_{0}$ is known. This is a common assumption in sequential changepoint literature.
3.1 Confidence set for $T$ with known pre-change distribution
[…]